Assume you have been asked to develop a master staffing plan for your facility. The facility...

  1. Assume you have been asked to develop a master staffing plan for your facility. The facility employees the following types of employees:
  1. Doctors
  2. Nurses
  3. Nurse Assistants
  4. Pharmacists
  5. Therapists
  6. Clerical

Assume the following:

  1. Your business year consists of 360 days
  2. Each employee works 5 rotating days. Each employee receives 2 days off per week
  3. Each employee receives 25 personal leave days and 10 floating holidays.

Calculate how many net paid days are worked and convert them into a Factor.
Answer #1

(i) Calculation of Net paid days worked by each employee:

Total no. of days worked in a week(a) = 5 days

No. Of weeks in a year(b) = 52 weeks

Total no. Of days worked in a year by each employee (x=a*b) = 260 days

Total no. of holidays receive by each employee (y) = 25 personal leave + 10 floating holidays = 35 days

Total net paid days worked by each employee (x-y)= 260 days - 35 days = 225 days

(ii) Factor = Total days in a business year/ Net paid days worked for facility

360 days/225 days = 1.6

Note: Net paid days worked are calculated for each employee. Total net paid days worked for facility can't be calculated as no. of staff is not given question.